Β π‘Šπ‘‹π‘Œπ‘ is a parallelogram with π‘Š(0, βˆ’3), 𝑋(5, βˆ’1), and π‘Œ(3, βˆ’4). Find the coordinates of point 𝑍. Recall that parallelogram points are listed in clockwise order.


1
Expert's answer
2021-04-14T08:11:36-0400

WXYZ

W(x,y) =(0,-3) ,X(x, y)=(5,-1) ,Y(x, y)=(3,-4) , Z(x, y)=(x1,y1)

parallelogram opposite sides are equal:

W-X=Y-Z AND W-Z=X-Y

W-Z= x x-xw and y=yx -yw , 5-0=5 , -1-(-3)=2

therefore difference is (5,2) implying:

Y-Z= xy-xz=5 and yy-yz=2

3-xz=5 and -4-yz =2

hence Z=(-2,-6)